The following data relating to units shipped and total shipping expense have been assembled by Archer Company, a manufacturer of large, custom-built air-conditioning units for commercial buildings:

Month Units
Shipped
Total Shipping
Expense
  January 10 $ 5,650        
  February 7 4,080        
  March 7 3,980    
  April 5 3,150        
  May 4 2,510        
  June 18 8,820      
  July 3 2,220        
Required:
1. Estimate a cost formula for shipping expense using the high-low method.
y= + x

      

2. Not available in Connect.
3.

What factors, other than the number of units shipped, are likely to affect the company’s total shipping expense? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer. Any boxes left with a question mark will be automatically graded as incorrect.)

    
Weight of the units shipped
Distance travelled
Number of workers used to manufacture the goods shipped
Size of the units shipped
Fixed costs to produce the units
Answer #1

Cost formula for shipping expense using the high-low method:-

It is computed using a few steps.

First we have to calculate fixed and variable costs.

Calculate variable cost per unit using identified high and low activity levels:

Variable cost

               = (Total cost of high activity- total cost of low activity) / (Highest activity unit- Lowest activity unit)

               =(8820-2220) / (18-3)

               =6600/15

               =440/ unit

Calculater fixed costs

To calculate the total fixed costs, use either the high or low cost and the variable cost into the total cost formula:

Here, we using the highest cost of june month

Total cost= (Variable cost*units produced)+ Total fixed cost

8820=(440*18)+Total fixed cost

Ie, Total fixed cost= 900

Using all these informations,

Total cost= Total fixed cost+ (Variable cost* units produced)

Total cost= 900+(440 * units produced)

Hence, cost formula for shipping expense using the high-low method:-

      Total cost= 900+(440 * units produced)

Ie,   y= 900 + 440 x

       Where y = total cost

                  x = units produced

(2)

What factors, other than the number of units shipped, are likely to affect the company’s total shipping expense?

Answer:

Weight of the units shipped: The shipping expense might increase if the weight of the shipment is very huge or very low.

Distance travelled

Generally shipping cost vary with the distance travelled